# Axis Tour: Word Tour Determines the Order of Axes in ICA-transformed Embeddings

### v2

@@ 6 Conclusion
In this study, we proposed a novel method, Axis Tour, which optimizes the order of axes in ICA-transformed word embeddings. We focus on the fact that the word embeddings reveal interpretable semantic axes while the order of these axes are arbitrary. Axis Tour aims to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. We also showed through experiments on downstream tasks that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
@@ Acknowledgements
We would like to thank Momose Oyama for the discussion. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110.
@@ A.1 Scatterplot drawing method
First, we define a set of axis indices for projection. Let $I:=\{a,\cdots,b\}\,(a,b\in[d],a<b)$ be a consecutive interval of indices. The number of indices in $I$ is $|I|=b-a+1$. For example, in Fig. 1, $a=86$ and $b=94$.
@@
Then we get the two-dimensional projection as $\mathbf{Q}_{I}:=\hat{\mathbf{T}}_{I}\mathbf{P}_{I}\in\mathbb{R}^{n\times 2}$. In $\mathbf{Q}_{I}$, the $\ell$th axis of $\hat{\mathbf{T}}$ is projected along the direction of $\varphi_{\ell}$.
We denote the $i$th word embedding of $\mathbf{Q}_{I}$ by $\mathbf{q}_{i}=(q_{i}^{x},q_{i}^{y})^{\top}\in\mathbb{R}^{2}$. When we plot the scatterplot of $\mathbf{Q}_{I}$, we do not plot the $i$th word embedding if $q_{i}^{y}<0$ for visual clarity, and show the top five words for each axis. The indices of these top words equal to the following index set $\text{Show}_{I}$ defined with $\text{Top}_{k}^{\ell}$ from section 4.1:
@@ A.2 Scatterplots of Table 1
(a) The 23rd axis to the 31st axis
(b) The 101st axis to the 109th axis
(c) The 237th axis to the 245th axis
@@
Fig. 5 shows the scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings in Table 1, using the procedure described in Appendix A.1. Similar to Fig. 1, it is evident that the top words of the axes of the Axis Tour embeddings are farther from the origin than those of the Skewness Sort, and the meanings of the adjacent axes change continuously.
@@ A.3 Evaluation metrics for scatterplots
In Section 1, we compared the quality of the scatterplots for Axis Tour and Skewness Sort by calculating the average distance of the top words from the origin. In this section, we first explain this metric and then, based on (4), define a new metric derived from the average of the cosine similarities between adjacent axis embeddings. We then compare these metrics for the scatterplots in Figs. 1 and 5. Similar to Appendix A.1, we use the Axis Tour embeddings to explain these metrics.
@@
Both $d_{I}$ and $c_{I}$ can also be calculated in a similar manner for Skewness Sort. Table 2 shows the values of $d_{I}$ and $c_{I}$ for Figs. 1 and 5. In these examples, Axis Tour shows higher values for both $d_{I}$ and $c_{I}$ compared to Skewness Sort, indicating better projection quality.
@@
Table 2: The values of $d_{I}$ and $c_{I}$ for Figs. 1 and 5.
@@ B.1.1 Skewness as weight
In (3), a vector $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for dimensionality reduction is defined by the skewness $\gamma_{\ell}\in\mathbb{R}_{\geq 0}$ of the $\ell$th axis of $\mathbf{T}$. In particular, $f_{r}^{(\ell)}$, which corresponds to the weight of the $\ell$th axis in the projection, is proportional to $\gamma_{\ell}^{\alpha}\,(\alpha\in\mathbb{R}_{\geq 0})$.
@@ B.1.2 Reason for Normalizing $\mathbf{f}_{r}$
Here, $\delta_{\ell r}$ is the Dirac delta function. From (13) the matrix $\mathbf{F}=[\mathbf{f}_{1},\cdots,\mathbf{f}_{d}]\in\mathbb{R}^{d\times d}$ corresponds to the $d$-dimensional identity matrix $\mathbf{I}\in\mathbb{R}^{d\times d}$, so $\mathbf{T}\mathbf{F}=\mathbf{T}\mathbf{I}=\mathbf{T}$.
@@ B.2 Projection from subspace to one-dimensional space
This section explains the projection from a subspace to a one-dimensional space using a specific example. Consider the subspace spanned by three consecutive axes (89, 90, 91) from Fig. 1. Fig. 6 shows the projection of this subspace using $\mathbf{f}_{r}$. The projection direction is in the direction representing the subspace, and the top words of each axis are projected close together.
@@ E.2 Comparison of $\alpha$
Fig. 7 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the PCA-transformed embeddings and the Axis Tour embeddings with $\alpha=0,1/3,1/2,1$.
@@ E.3.1 Selection of $k$
Similar to $\hat{\mathbf{S}}$, we define the matrix $\hat{\mathbf{T}}_{k_{1}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{T}_{k_{1}}$ with row vectors $\hat{\mathbf{t}}_{k_{1},i}=\mathbf{t}_{k_{1},i}/\|\mathbf{t}_{k_{1},i}\|$. Here, the $i$th word embedding of $\mathbf{T}_{k_{1}}$ and $\hat{\mathbf{T}}_{k_{1}}$ are denoted by $\mathbf{t}_{k_{1},i},\hat{\mathbf{t}}_{k_{1},i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$th axis of $\hat{\mathbf{T}}_{k_{1}}$ and denote the index set of words corresponding to the top $k_{2}$ elements as $\text{Top}_{k_{2}}^{\ell}$.
We redefine the $\ell$th axis embedding $\mathbf{v}_{\ell}{(k_{1},k_{2})}$ for $k_{1}$ and $k_{2}$ as follows:
@@
For example, $C_{1}(k)$ in (18) is a function that measures the semantic continuity for the axis embedding $\mathbf{v}_{\ell}{(1,k)}$ redefined by the top $k$ words of the $\ell$th axis of $\mathbf{T}_{1}$. To evaluate the robustness of $\mathbf{T}_{1}$ to changes for $k$, we average the values of $C_{1}(k)$ for $k=1,10,100,1000$. A higher average value indicates better quality.
@@ Results.
Fig. 8 shows plots of the functions $C_{1}(k)$, $C_{10}(k)$, $C_{100}(k)$, $C_{1000}(k)$ and $M(k)$ for $k=1,10,100,1000$. $M(100)$ is the maximum value of $M(k)$. This result validates the use of $k=100$ as the default value for our experiment settings.
E.3.2 Qualitative observation
@@ E.3.3 Dimensionality reduction
Fig. 9 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the PCA-transformed embeddings and the Axis Tour embeddings with $k=1,10,100,1000$ and $\alpha=1/3$.
@@ E.4 Dimensionality reduction by projection for Skewness Sort and Rand Order
For the sake of explanation, let the embedding matrices for Skewness Sort and Random Order be denoted as $\mathbf{S}_{\text{skew}}$ and $\mathbf{S}_{\text{rand}}$, respectively. Let the skewness of the $\ell$-th axis for each be $\gamma_{\text{skew},\ell},\gamma_{\text{rand},\ell}\in\mathbb{R}$. By the definition of Skewness Sort, $\gamma_{\text{skew},\ell}\geq 0$. Similar to Section 4.3, we consider reducing the dimensions from $d$ to $p(\leq d)$, divide $[d]$ into $p$ equal-length intervals, and use the index set for the $r$th interval $I_{r}=\{a_{r},\cdots,b_{r}\}\,(a_{r},b_{r}\in[d],a_{r}\leq b_{r})$.
@@ E.4.1 Skewness Sort Projection
where $\alpha\in\mathbb{R}_{\geq 0}$. We then define $\mathbf{F}_{\text{skew}}:=[\mathbf{f}_{\text{skew},1},\cdots,\mathbf{f}_{\text{skew},p}]\in\mathbb{R}^{d\times p}$ and obtain the $p$-dimensional embedding matrix $\mathbf{S}_{\text{skew}}\mathbf{F}_{\text{skew}}\in\mathbb{R}^{n\times p}$. We call these embeddings as Skewness Sort Projection.
@@ E.4.2 Random Order Projection
where $\alpha\in\mathbb{R}_{\geq 0}$ and $\text{sgn}(\cdot)$ represents the sign function. We then define $\mathbf{F}_{\text{rand}}:=[\mathbf{f}_{\text{rand},1},\cdots,\mathbf{f}_{\text{rand},p}]\in\mathbb{R}^{d\times p}$ and obtain the $p$-dimensional embedding matrix $\mathbf{S}_{\text{rand}}\mathbf{F}_{\text{rand}}\in\mathbb{R}^{n\times p}$. We call these embeddings as Random Order Projection.
@@ E.4.3 Results
Similar to Axis Tour, we set the default value of $\alpha$ to $\alpha=1/3$ for both Skewness Sort Projection and Random Order Projection. Fig. 10 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the embeddings.
@@ E.5.1 Qualitative observation
Table 9: The values of $d_{I}$ and $c_{I}$ for Figs. 11 and 12.
@@ E.5.2 Scatterplots of Table 8
Figs. 11 and 12 show the scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings for word2vec and BERT in Tables 8(a) and 8(b), respectively. We used the procedure described in Appendix A.1. Similar to Figs. 1 and 5, for both word2vec and BERT, we can see that the top words of the axes are farther from the origin than those of the Skewness Sort, and the meanings of the adjacent axes change continuously.
For Figs. 11 and 12, Table 9 shows the values of $d_{I}$ and $c_{I}$ as the evaluation metrics defined in Appendix A.3. Similar to Table 2, for both word2vec and BERT, Axis Tour shows higher values for both $d_{I}$ and $c_{I}$ than Skewness Sort, indicating better projection quality.
@@ E.5.3 Cosine similarity between adjacent axis embeddings
Fig. 13 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ of the Axis Tour embeddings and the baseline embeddings for both word2vec and BERT. Similar to the results for GloVe in Fig. 2, the values of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ are consistently higher in Axis Tour, while this trend is not observed in the other baselines.
Figs. 14 and 15 illustrate the relationship between the skewness and the average of consecutive two cosine similarities for Axis Tour and Skewness Sort. Interestingly, while the Axis Tour embeddings for BERT show a strong correlation between the skewness and the average of two cosine similarities in Fig. 15(a), similar to the results for GloVe in Fig. 3(a), the Axis Tour embeddings for word2vec do not show such a correlation in Fig. 14(a).
@@ E.5.4 Dimensionality reduction: analogy, word similarity, and categorization tasks
Fig. 16 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the embeddings derived from word2vec. Axis Tour outperformed both Random Order and Skewness Sort in each task. In analogy tasks, Axis Tour outperformed PCA in lower dimensions and was nearly equivalent in other dimensions.
@@
Appendix F Topographic ICA (TICA)

### v3

@@ 7 Conclusion
In this study, we proposed a novel method, Axis Tour, which optimizes the order of axes in ICA-transformed word embeddings. We focused on the fact that the word embeddings reveal interpretable semantic axes while the order of these axes is arbitrary. Axis Tour aims to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. Additionally, we demonstrated through experiments on downstream tasks that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
@@ Acknowledgements
We would like to thank Momose Oyama for the discussion and anonymous reviewers for their helpful advice. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110.
Code availability
Our code is available at https://github.com/ymgw55/Axis-Tour.
@@ A.1 Scatterplot drawing method
First, we define a set of axis indices for projection. Let $I:=\{a,\ldots,b\}\,(a,b\in[d],a<b)$ be a consecutive interval of indices. The number of indices in $I$ is $|I|=b-a+1$. For example, in Fig. 1, $a=86$ and $b=94$.
@@
Then we get the two-dimensional projection as $\mathbf{Q}_{I}:=\hat{\mathbf{T}}_{I}\mathbf{P}_{I}\in\mathbb{R}^{n\times 2}$. In $\mathbf{Q}_{I}$, the $\ell$-th axis of $\hat{\mathbf{T}}$ is projected along the direction of $\varphi_{\ell}$.
We denote the $i$-th word embedding of $\mathbf{Q}_{I}$ by $\mathbf{q}_{i}=(q_{i}^{x},q_{i}^{y})^{\top}\in\mathbb{R}^{2}$. When we plot the scatterplot of $\mathbf{Q}_{I}$, we do not plot the $i$-th word embedding if $q_{i}^{y}<0$ for visual clarity, and show the top five words for each axis. The indices of these top words equal to the following index set $\text{Show}_{I}$ defined with $\text{Top}_{k}^{\ell}$ from section 4.1:
@@ A.2 Scatterplots of Table 1
(c) The 23rd axis to the 31st axis
(d) The 101st axis to the 109th axis
(e) The 237th axis to the 245th axis
@@
Figure 6 shows the scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings in Table 1, using the procedure described in Appendix A.1. Similar to Fig. 1, it is evident that the top words of the axes of the Axis Tour embeddings are farther from the origin than those of the Skewness Sort, and the meanings of the adjacent axes change continuously.
@@ A.3 Evaluation metrics for scatterplots
In Section 1, we compared the quality of the scatterplots for Axis Tour and Skewness Sort by calculating the average distance of the top words from the origin. In this section, we first explain this metric and then, based on (4), define a new metric derived from the average of the cosine similarities between adjacent axis embeddings. We then compare these metrics for the scatterplots in Figs. 1 and 6. Similar to Appendix A.1, we use the Axis Tour embeddings to explain these metrics.
@@
Both $d_{I}$ and $c_{I}$ can also be calculated in a similar manner for Skewness Sort. Table 2 shows the values of $d_{I}$ and $c_{I}$ for Figs. 1 and 6. In these examples, Axis Tour shows higher values for both $d_{I}$ and $c_{I}$ compared to Skewness Sort, indicating better projection quality.
@@
Table 2: The values of $d_{I}$ and $c_{I}$ for Figs. 1 and 6.
@@ B.1.1 Skewness as weight
In (3), a vector $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for dimensionality reduction is defined by the skewness $\gamma_{\ell}\in\mathbb{R}_{\geq 0}$ of the $\ell$-th axis of $\mathbf{T}$. In particular, $f_{r}^{(\ell)}$, which corresponds to the weight of the $\ell$-th axis in the projection, is proportional to $\gamma_{\ell}^{\alpha}\,(\alpha\in\mathbb{R}_{\geq 0})$.
@@ B.1.2 Reason for Normalizing $\mathbf{f}_{r}$
Here, $\delta_{\ell r}$ is the Dirac delta function. From (13) the matrix $\mathbf{F}=[\mathbf{f}_{1},\ldots,\mathbf{f}_{d}]\in\mathbb{R}^{d\times d}$ corresponds to the $d$-dimensional identity matrix $\mathbf{I}\in\mathbb{R}^{d\times d}$, so $\mathbf{T}\mathbf{F}=\mathbf{T}\mathbf{I}=\mathbf{T}$.
@@ B.2 Projection from subspace to one-dimensional space
This section explains the projection from a subspace to a one-dimensional space using a specific example. Consider the subspace spanned by three consecutive axes (89, 90, 91) from Fig. 1. Figure 7 shows the projection of this subspace using $\mathbf{f}_{r}$. The projection direction is in the direction representing the subspace, and the top words of each axis are projected close together.
@@ E.2 Comparison of $\alpha$
Figure 8 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the PCA-transformed embeddings and the Axis Tour embeddings with $\alpha=0,1/3,1/2,1$.
@@ E.3.1 Selection of $k$
Similar to $\hat{\mathbf{S}}$, we define the matrix $\hat{\mathbf{T}}_{k_{1}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{T}_{k_{1}}$ with row vectors $\hat{\mathbf{t}}_{k_{1},i}=\mathbf{t}_{k_{1},i}/\|\mathbf{t}_{k_{1},i}\|$. Here, the $i$-th word embedding of $\mathbf{T}_{k_{1}}$ and $\hat{\mathbf{T}}_{k_{1}}$ are denoted by $\mathbf{t}_{k_{1},i},\hat{\mathbf{t}}_{k_{1},i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$-th axis of $\hat{\mathbf{T}}_{k_{1}}$ and denote the index set of words corresponding to the top $k_{2}$ elements as $\text{Top}_{k_{2}}^{\ell}$.
We redefine the $\ell$-th axis embedding $\mathbf{v}_{\ell}{(k_{1},k_{2})}$ for $k_{1}$ and $k_{2}$ as follows:
@@
For example, $C_{1}(k)$ in (18) is a function that measures the semantic continuity for the axis embedding $\mathbf{v}_{\ell}{(1,k)}$ redefined by the top $k$ words of the $\ell$-th axis of $\mathbf{T}_{1}$. To evaluate the robustness of $\mathbf{T}_{1}$ to changes for $k$, we average the values of $C_{1}(k)$ for $k=1,10,100,1000$. A higher average value indicates better quality.
@@ Results.
Figure 9 shows plots of the functions $C_{1}(k)$, $C_{10}(k)$, $C_{100}(k)$, $C_{1000}(k)$ and $M(k)$ for $k=1,10,100,1000$. $M(100)$ is the maximum value of $M(k)$. This result validates the use of $k=100$ as the default value for our experiment settings.
E.3.2 Qualitative observation
@@ E.3.3 Dimensionality reduction
Figure 10 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the PCA-transformed embeddings and the Axis Tour embeddings with $k=1,10,100,1000$ and $\alpha=1/3$.
@@ E.4 Dimensionality reduction by projection for Skewness Sort and Rand Order
For the sake of explanation, let the embedding matrices for Skewness Sort and Random Order be denoted as $\mathbf{S}_{\text{skew}}$ and $\mathbf{S}_{\text{rand}}$, respectively. Let the skewness of the $\ell$-th axis for each be $\gamma_{\text{skew},\ell},\gamma_{\text{rand},\ell}\in\mathbb{R}$. By the definition of Skewness Sort, $\gamma_{\text{skew},\ell}\geq 0$. Similar to Section 4.3, we consider reducing the dimensions from $d$ to $p(\leq d)$, divide $[d]$ into $p$ equal-length intervals, and use the index set for the $r$-th interval $I_{r}=\{a_{r},\ldots,b_{r}\}\,(a_{r},b_{r}\in[d],a_{r}\leq b_{r})$.
@@ E.4.1 Skewness Sort Projection
where $\alpha\in\mathbb{R}_{\geq 0}$. We then define $\mathbf{F}_{\text{skew}}:=[\mathbf{f}_{\text{skew},1},\ldots,\mathbf{f}_{\text{skew},p}]\in\mathbb{R}^{d\times p}$ and obtain the $p$-dimensional embedding matrix $\mathbf{S}_{\text{skew}}\mathbf{F}_{\text{skew}}\in\mathbb{R}^{n\times p}$. We call these embeddings as Skewness Sort Projection.
@@ E.4.2 Random Order Projection
where $\alpha\in\mathbb{R}_{\geq 0}$ and $\text{sgn}(\cdot)$ represents the sign function. We then define $\mathbf{F}_{\text{rand}}:=[\mathbf{f}_{\text{rand},1},\ldots,\mathbf{f}_{\text{rand},p}]\in\mathbb{R}^{d\times p}$ and obtain the $p$-dimensional embedding matrix $\mathbf{S}_{\text{rand}}\mathbf{F}_{\text{rand}}\in\mathbb{R}^{n\times p}$. We call these embeddings as Random Order Projection.
@@ E.4.3 Results
Similar to Axis Tour, we set the default value of $\alpha$ to $\alpha=1/3$ for both Skewness Sort Projection and Random Order Projection. Figure 11 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the embeddings.
@@ E.4.4 Setting of $p=d\,(=300)$
Table 9: The values of $d_{I}$ and $c_{I}$ for Figs. 12 and 13.
E.5 Other embeddings
E.5.1 Qualitative observation
@@ E.5.2 Scatterplots of Table 8
Figures 12 and 13 show the scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings for word2vec and BERT in Tables 8(a) and 8(b), respectively. We used the procedure described in Appendix A.1. Similar to Figs. 1 and 6, for both word2vec and BERT, we can see that the top words of the axes are farther from the origin than those of the Skewness Sort, and the meanings of the adjacent axes change continuously.
For Figs. 12 and 13, Table 9 shows the values of $d_{I}$ and $c_{I}$ as the evaluation metrics defined in Appendix A.3. Similar to Table 2, for both word2vec and BERT, Axis Tour shows higher values for both $d_{I}$ and $c_{I}$ than Skewness Sort, indicating better projection quality.
@@ E.5.3 Cosine similarity between adjacent axis embeddings
Figure 14 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ of the Axis Tour embeddings and the baseline embeddings for both word2vec and BERT. Similar to the results for GloVe in Fig. 2, the values of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ are consistently higher in Axis Tour, while this trend is not observed in the other baselines.
Figures 15 and 16 illustrate the relationship between the skewness and the average of consecutive two cosine similarities for Axis Tour and Skewness Sort. Interestingly, while the Axis Tour embeddings for BERT show a strong correlation between the skewness and the average of two cosine similarities in Fig. 16(a), similar to the results for GloVe in Fig. 5(a), the Axis Tour embeddings for word2vec do not show such a correlation in Fig. 15(a).
@@ E.5.4 Dimensionality reduction: analogy, word similarity, and categorization tasks
Figure 17 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the embeddings derived from word2vec. Axis Tour outperformed both Random Order and Skewness Sort in each task. In analogy tasks, Axis Tour outperformed PCA in lower dimensions and was nearly equivalent in other dimensions.
@@ Appendix F Topographic ICA (TICA)
F.1 Overview of TICA
$\sigma_{\ell}$ $=\phi\left(\sum_{m=1}^{d}h_{\ell m}u_{m}\right),$ (25)
$s_{\ell}$ $=\sigma_{\ell}z_{\ell},$ (26)
where $\phi$ is some nonlinear function, $h_{\ell m}$ is a symmetric neighborhood relation, and $z_{\ell}$ are mutually independent random variables. To estimate the decomposition matrix $\mathbf{W}=[\mathbf{w}_{1},\ldots,\mathbf{w}_{d}]^{\top}\in\mathbb{R}^{d\times d}$, where $\mathbf{w}_{\ell}\in\mathbb{R}^{d}$, such that $\mathbf{s}=\mathbf{W}\mathbf{x}$ from the mixed signals $\mathbf{x}\in\mathbb{R}^{d}$, TICA maximize the log likelihood
The optimization of (28) can be done by gradient descent:
Note that we have to apply orthonormalization
$\mathbf{W}\leftarrow(\mathbf{WW}^{\top})^{-1/2}\mathbf{W}$ (30)
$\mathbf{S}=\mathbf{X}\mathbf{W}^{\top}.$ (31)
F.2 Experiments
As in classic ICA, we transform the $300$-dimensional GloVe by TICA for our experiments. Following the setup of the 1-D experiment in Hyvärinen et al. (2001a), a neighborhood relation was defined by convolving a rectangular-shaped filter $(\ldots,0,0,0,\overbrace{1,1,\ldots,1,1}^{\mathsf{width}~{}\text{times}},0,0,0,\ldots)$ twice with itself. Hyvärinen et al. (2001a) defined the hyperparameter $\mathsf{width}=5$ for 20 components, so for 300 components, we consider two cases where the width is scaled logarithmically or linearly to $\mathsf{width}=9$ and $\mathsf{width}=75$. These are denoted as TICA9 and TICA75, respectively [^22] For the TICA9-transformed and TICA75-transformed embeddings, similar to the Axis Tour embeddings, we flip the sign of each axis as needed so that the skewness is positive.. We performed $10{,}000$ iterations of gradient descent in both settings.
F.2.1 Qualitative observation